You are the veterinarian at a zoo that has 21 zebras. Each zebra needs to get four intravenous shots (over the course of a month) of a vaccination against Zebra Zombie Disease. Each injection contains 750 mg of the active vaccination agent. How many grams of the active vaccination agent should you order for the month?
1
Expert's answer
2020-09-16T10:12:36-0400

To vaccinate one zebra during the month the doctor needs to order (assume 30 days in month) "750\dfrac{mg}{day}\times 30days = 22500mg = 22.5g" of the active vaccination agent. To vaccinate all 21 zebras he should order "22.5g\times 21 = 472.5g" of the active vaccination agent.


Answer. 472.5 g.